(S-C)/C = R


I am trying to solve for C. Can you rearrage the formula so C is on the right side:
ex. some+formula = C
Thanks!!

Answer by checkley77(12844)

(S-C)/C=R CROSS MULTIPLY.
S-C=CR ADD C TO BOTH SIDES.
S=CR+C FACTOR OUT C.
S=C(R+1) DIVIDE BY (R+1)
S/(R+1)=C ANS.